# Breaking the Curse of Multilinguality with Cross-lingual Expert Language Models

### v1

@@ Abstract
Despite their popularity in non-English NLP, multilingual language models often underperform monolingual ones due to inter-language competition for model parameters. We propose Cross-lingual Expert Language Models (x-elm), which mitigate this competition by independently training language models on subsets of the multilingual corpus. This process specializes x-elms to different languages while remaining effective as a multilingual ensemble. Our experiments show that when given the same compute budget, x-elm outperforms jointly trained multilingual models across all considered languages and that these gains transfer to downstream tasks. x-elm provides additional benefits over performance improvements: new experts can be iteratively added, adapting x-elm to new languages without catastrophic forgetting. Furthermore, training is asynchronous, reducing the hardware requirements for multilingual training and democratizing multilingual modeling.
$\dagger$\$$\dagger$\$footnotetext: Correspondence to blvns@cs.washington.edu**footnotetext: Work done while visiting the University of Washington
@@ 1 Introduction
Massively multilingual language models (LMs), which are trained on terabytes of text in a hundred or more languages, underlie almost all non-English and cross-lingual NLP applications (Scao et al., 2022; Lin et al., 2022, inter alia). Despite their wide adoption, these models come at a cost: by modeling many languages in a single model, there is inter-language competition for fixed model capacity; this causes performance on individual languages to degrade relative to monolingual models Conneau et al. (2020); Chang et al. (2023). Furthermore, this phenomenon (termed the curse of multilinguality) can significantly harm low-resource languages Wu and Dredze (2020).
@@
In this paper, we address the curse of multilinguality with Cross-lingual Expert Language Models (x-elm, Figure 1), an ensemble of language models initialized from a pretrained multilingual model and each independently trained on a different subset of a multilingual corpus. Our ensemble allows for efficient scaling of model capacity to better represent all the corpus languages. These x-elms are trained with x-BTM, a new extension of the Branch-Train-Merge paradigm (BTM; Li et al., 2022; Gururangan et al., 2023, §2) to the more heterogenous multilingual setting.
x-BTM improves over existing BTM techniques by introducing (1) a new method for balanced clustering of multilingual data based on typological similarity and (2) Hierarchical Multi-Round training (HMR), an algorithm for efficiently training new experts specialized to unseen languages or other distributions of multilingual data. Once the initial x-elms are trained, we dynamically select experts to perform inference (§3.3). We can also efficiently adapt x-elms to novel settings with additional rounds of x-BTM on new experts branched from existing x-elms (§4); this improves the overall x-elm set without altering the existing experts.
@@
We train x-elms on 20 total languages—including adapting to 4 unseen ones–and on up to 21 billion training tokens. Our experiments demonstrate that x-elms outperform the dense language models given the same compute budget in every considered experimental setting, with improvements of up to 3.8 perplexity points (§6). Furthermore, the perplexity gains observed in x-elm languages are well-balanced across language resourcedness, and adapting the models to new languages via HMR training significantly outperforms standard language-adaptive pretraining methods. We also show that the language modeling gains of x-elm hold on downstream task evaluations (§7).
@@
Multilingual modeling with x-elm provides additional benefits over improved performance. Training a set of x-elms is more computationally efficient than a comparable dense model; each expert is trained independently, which removes the overhead cost of cross-GPU synchronization Li et al. (2022) and allows experts to be trained asynchronously in low-compute settings. Similarly, adapting x-elms to new languages is more efficient than continued training of a dense LM and does not risk catastrophic forgetting of previously seen languages, as adding a new x-elm does not change the existing experts. As a result, x-elms allow much more efficient modeling than prior multilingual approaches, democratizing work on building and improving multilingual systems.
@@ 2 Background: Branch-Train-Merge
Branch-Train-Merge (BTM; Li et al. 2022) alleviates this cost by dividing the total compute among smaller expert language models that are trained independently on different domains (or subsets of a corpus) and then combined during inference time. While the total number of parameters increases with the number of experts, inference with these models often uses a subset of experts (see §3.3), keeping inference costs manageable.
c-BTM Gururangan et al. (2023) generalizes the above approach with cluster-based representations of domains. Across multiple corpora, they show that (1) the optimal number of experts increases with data and compute and (2) a set of small expert models performs similarly to equivalently sized dense models at vastly reduced FLOP budgets.
Our work extends these studies to the multilingual setting, in which experts are specialized to different languages instead of (primarily) English-language domains. In the multilingual setting, we can also use typological structure to specialize experts, which we show provides additional benefits over automatic data clustering. We also demonstrate that training along the hierarchy of language families in multiple rounds yields further performance benefits.
@@ 3 Cross-lingual Expert Language Models
Multilingual language models are jointly trained on many different languages (e.g., Lin et al., 2022), despite the well-documented curse of multilinguality that comes from the competition between languages for fixed model capacity Conneau et al. (2020); Wang et al. (2020). We propose Cross-lingual Expert Language Models, or x-elms, to address this performance disparity (Figure 1). These experts are trained with x-BTM, an extension of the Branch-Train-Merge (BTM) pretraining paradigm Li et al. (2022); Gururangan et al. (2023): we asynchronously train many expert LMs on subsets of a multilingual corpus in order to specialize them to different sub-distributions of the multilingual space and then merge the experts to perform inference. We hypothesize that this training scheme will alleviate the curse of multilinguality on individual languages while maintaining the cross-lingual properties of dense multilingual LMs.
@@
This section overviews our algorithm for sparse training of multilingual experts.
@@
As a preprocessing step, we partition the multilingual corpus into $k$ clusters to train each x-elm. We consider learning TF-IDF clusters as well as a new clustering method that groups documents by language identity and linguistic typology (§3.2).
@@
A preliminary stage of shared, dense pretraining is important for ensembling expert language models Li et al. (2022). Therefore, the first step of BTM is to initialize (branch) each expert with the parameters from a partially trained model. For this work, we initialize our x-elms with an existing multilingual pretrained model, XGLM Lin et al. (2022).
@@
Steps 1 – 3 describe a single round of x-BTM training. However, we can continue to update the x-elm set by branching—initializing a new group of experts—from existing models in the ensemble and performing more rounds of x-BTM via the method we propose in §4. This allows us to further improve x-ELM by training and adding new experts.
@@ 3.2 Data Allocation Methods
How we assign data to experts is a key component of training x-elm, and it is a particularly crucial choice as the data becomes more diverse (i.e., spanning many languages). We consider two methods of data allocation when training our x-elms:
@@
We partition the multilingual corpus automatically into $k$ components with k-means clustering. First, we encode each document into a word-level TF-IDF representation [^2] Data tokenization is independent of the downstream model. Here, we use the sklearn text-vectorizer tokenizer.; we then perform balanced k-means clustering on these representations to obtain approximately balanced subsets of the data on which to train each x-elm. Further details on the balanced k-means clustering method can be found in Gururangan et al. (2023). This allocation method uses no language information outside of what is inherent in the text (e.g., script, vocabulary).
Figure 2: Heirachical clustering of languages used to train our x-elm ensembles.
@@
We also consider segmenting the corpus by language identity. [^3] This requires knowledge of the language of each document. We use the language tags provided with mC4. Rather than balancing the amount of data allocated to each cluster in this setting, we instead keep the number of languages per cluster fixed. Specifically, we learn a balanced hierarchical clustering of the languages (Figure 2). We build this hierarchy using the language similarity metrics in lang2vec Littell et al. (2017), which represents languages based on linguistic features in resources such as WALS [^4] World Atlas of Language Structures, https://wals.info/ and estimates language similarity with distance in this feature space. We first initialize each cluster with a single language; at each step, we merge each cluster with exactly one other based on the minimum distances between the cluster centroids. We then group languages according to the resulting hierarchy and the desired number of experts. When the number of languages equals the number of experts, typological clustering results in monolingual training, where every language is assigned a separate expert.
Comparing the Clustering Techniques
@@ 3.3 Inference with x-elms
We evaluate a number of different methods for performing inference with x-elms:
@@
This method performs inference with a single expert chosen prior to evaluation; therefore, it incurs the same inference cost as the dense baselines. When evaluating the Typology experts on a particular language $\ell$, we choose the expert that included $\ell$ in the set of languages on which they continued pretraining. Similarly, when evaluating TF-IDF, we choose the x-elm trained on the highest percentage of $\ell$’s data.
@@
We also consider ensembling TF-IDF experts by adapting the c-BTM ensemble routing method. Here, we calculate ensembling alphas, or weights, over these experts for each evaluation step based on the proceeding context’s TF-IDF distance from the experts’ k-means centroids. These weights are then used to ensemble the output probabilities from each expert.
More specifically, given a probability from each expert LM $p_{e}(x_{t}|x_{<t})$ and the corresponding ensemble weight $\alpha_{e}=p(e|x_{<t})\propto\textrm{exp}(-\textrm{dist}(x_{<t},c_{e})^{2}/T)$, the probability of the ensemble $p_{E}(x_{t}|x_{<t})=\sum_{e\in E}\alpha_{e}\cdot p_{e}(x_{t}|x_{<t})$. Here, $\textrm{dist}(x_{<t},c_{e})$ is obtained by embedding $x_{<t}$ with the learned TF-IDF vectorizer and calculating the Euclidean distance from $c_{e}$ (the centroid over the data representations allocated to expert $e$), and $T$ is a temperature parameter over the ensemble weight distribution. Further details and motivation for this setting are provided in Gururangan et al. (2023).
Ensembling x-elm outputs increases the cost of inference relative to the dense model or top-1 inference. However, it can potentially better fit different subsets of data in a diverse evaluation set. We also do not assume we know the identity of each example when ensembling, which makes this approach more flexible than the top-1 setting. In most cases, we ensemble all $k$ experts; however, we can also reduce computational costs by sparsifying the ensemble weights and only activating the $m$ ($<k$) experts that most contribute to an example: $p_{E}(x_{t}|x_{<t})=\sum_{e\in E}\alpha_{e}\cdot p_{e}(x_{t}|x_{<t}):\alpha_{e}\in\textrm{top-m}(\bf{\alpha}_{E})$. Table 8 presents the performance tradeoff with sparser TF-IDF ensembles.
@@ 4 Hierarchical Multi-Round Training
We previously described a single round of training for x-elm3.1). However, BTM can also be used repeatedly to train new experts seeded with those learned in a prior round. The multilingual setting provides a natural extension of multi-round training that leverages typological structure when initializing new experts.
@@
HMR pretraining gives multiple benefits over single-round BTM. In particular, HMR training saves compute and more easily adapts our x-elms to new settings. A specific application of this is adding new languages to the model: while updating dense multilingual LMs with new languages is difficult and can lead to catastrophic forgetting of existing languages Winata et al. (2023), hierarchically training an expert on a new language adds it to the x-elm set without altering the existing information in other experts. We further consider this use case for HMR training in §6.3.
[…]
@@ 5 Experimental Design
We present a series of experiments to test whether the x-elm pretraining paradigm remedies the decrease in individual language performance observed in dense multilingual models.
@@ 5.1 Pretraining Data and Languages
We train our x-elms on mC4, an open-source, multilingual pretraining corpus derived from CommonCrawl Xue et al. (2021). [^5] While one could also continue pretraining with the same corpus that the seed LM was trained on, the pretraining data for XGLM is not publicly available. mC4 provides language tags for each document in the corpus, which were automatically assigned with cld3666https://github.com/google/cld3 when the dataset was constructed; we use these language tags during typological clustering (§3.2). We focus our experiments on the 16 highest-resourced languages out of the 30 languages on which the seed LM, XGLM-1.7B, was trained. For languages with significantly more data than the others (e.g., English), we subsample their data to the first 1,024 shards. Appendix Table 5 gives the languages and data quantities in our pretraining corpus.

### v2

@@ Abstract
Despite their popularity in non-English NLP, multilingual language models often underperform monolingual ones due to inter-language competition for model parameters. We propose Cross-lingual Expert Language Models (x-elm), which mitigate this competition by independently training language models on subsets of the multilingual corpus. This process specializes x-elms to different languages while remaining effective as a multilingual ensemble. Our experiments show that when given the same compute budget, x-elm outperforms jointly trained multilingual models across all 16 considered languages and that these gains transfer to downstream tasks. x-elm provides additional benefits over performance improvements: new experts can be iteratively added, adapting x-elm to new languages without catastrophic forgetting. Furthermore, training is asynchronous, reducing the hardware requirements for multilingual training and democratizing multilingual modeling.
Breaking the Curse of Multilinguality with Cross-lingual
Expert Language Models
Terra Blevins1†  Tomasz Limisiewicz2∗  Suchin Gururangan1  Margaret Li1 Hila Gonen1  Noah A. Smith1,3  Luke Zettlemoyer1 1Paul G. Allen School of Computer Science and Engineering, University of Washington 2Faculty of Mathematics and Physics, Charles University in Prague 3Allen Institute for Artificial Intelligence
$\dagger$\$$\dagger$\$footnotetext: Correspondence to blvns@cs.washington.edu**footnotetext: Work done while visiting the University of Washington.
@@ 1 Introduction
Massively multilingual language models (LMs), which are trained on terabytes of text in a hundred or more languages, underlie almost all non-English and cross-lingual NLP applications (Scao et al., 2022; Lin et al., 2022, i.a.). Despite their wide adoption, these models come at a cost: the many languages are represented in the same, fixed model capacity, causing performance on individual languages to degrade relative to monolingual models Conneau et al. (2020); Chang et al. (2023). This phenomenon (termed the curse of multilinguality) can significantly harm low-resource language performance Wu and Dredze (2020).
@@
In this paper, we address this curse with Cross-lingual Expert Language Models (x-elm, Figure 1), an ensemble of language models initialized from a pretrained multilingual model and each independently trained on a different portion of a multilingual corpus with x-BTM, a new extension of the Branch-Train-Merge paradigm (BTM; Li et al., 2022; Gururangan et al., 2023, §2) to the more heterogenous multilingual setting. x-elm allows for efficient scaling of model capacity to better represent all considered languages.
x-BTM adapts existing BTM techniques to the multilingual setting by introducing a new cluster method for data assignments based on typological language similarity (§3.2). We also propose Heirachical Multi-round Training (HMR; §4), a method for efficiently adapting trained x-elms to novel multilingual settings by branching from existing, typologically related x-elms; this method for adapting x-elm to new languages strongly outperforms standard language adaptation methods.
@@
We train x-elms on 20 languages—including adapting to 4 unseen ones—with up to 21 billion training tokens, with the 1.7B parameter XGLM model as the base Lin et al. (2022). Our experiments show that x-elm strongly outperforms the dense multilingual models given the same compute budget in every considered setting (§6), and that these improvements consistently benefit every language. Notably, monolingual experts generally underperform typologically-informed multilingual x-elms, indicating that linguistically targeted multilinguality can benefit language modeling. We then show that the language modeling gains of x-elm hold on downstream evaluations (§7).
@@
Multilingual modeling with x-elm provides additional benefits beyond improved performance. Training a set of x-elms is more computationally efficient than training a comparable dense model; each expert is trained independently, which removes the overhead cost of cross-GPU synchronization Li et al. (2022) and allows asynchronous model training in low-compute settings. Similarly, adapting x-elms to new languages with HMR training—a popular use case of language models (i.e., Chau et al., 2020)—is more efficient than continued training of a dense LM and does not risk forgetting previously learned languages Yogatama et al. (2019). As a result, x-elm allows much more efficient and effective multilingual modeling than prior approaches, democratizing multilingual NLP. We release the code and trained models. [^1] https://github.com/blvns/x-elm/
@@ 2 Background: Branch-Train-Merge
Branch-Train-Merge (BTM; Li et al. 2022) alleviates this cost by dividing the total compute among smaller expert language models that are trained independently on different domains and then combined during inference time. While the total number of parameters increases with the number of experts, inference with these models often uses a subset of experts (see §3.3), keeping inference costs manageable. c-BTM Gururangan et al. (2023) then generalizes the above approach with automatic clustering of domains. Across multiple corpora, they show that (1) the optimal number of experts increases with data and compute and (2) a set of small expert models performs similarly to equivalently sized dense models at vastly reduced FLOP budgets.
Our work extends these studies to the multilingual setting, where experts are specialized to different languages instead of English-language domains. In the multilingual setting, we can also use typological structure to specialize experts, which we show provides additional benefits over automatic clustering. We also demonstrate that leveraging the hierarchy of language families in multi-round training yields further performance gains.
@@ 3 Cross-lingual Expert Language Models
Multilingual language models are jointly trained on many different languages (e.g., Lin et al., 2022), despite the well-documented effect this has on individual language performance Conneau et al. (2020); Wang et al. (2020). We propose Cross-lingual Expert Language Models, or x-elms (Figure 1), which we hypothesize will alleviate the curse of multilinguality while maintaining the cross-lingual properties of dense multilingual LMs.
@@
This section overviews our algorithm for the sparse training of multilingual experts.
@@
As a preprocessing step, we partition the multilingual corpus into $k$ clusters to train each x-elm. We consider both TF-IDF clusters and a new clustering method that groups documents by language identity and linguistic typology (§3.2).
@@
A preliminary stage of shared, dense pretraining is important for ensembling expert language models Li et al. (2022). Therefore, the first step of BTM is to initialize (branch) each expert with the parameters from a partially trained language model. For this work, we initialize our x-elms with an existing multilingual pretrained model, XGLM Lin et al. (2022).
@@
Steps 1–3 describe a single round of x-BTM training. However, we can continue to update the x-elm set by branching—initializing a new group of experts—from existing models in the ensemble and performing more rounds of x-BTM via the method proposed in §4. This allows us to further improve x-ELM by training and adding new experts.
@@ 3.2 Data Allocation Methods
How we assign data to experts is a key component of training x-elm, particularly as the data becomes more diverse (i.e., spanning many languages). We consider two methods of data allocation:
@@
We partition the multilingual corpus automatically into $k$ components with $k$-means clustering. First, we encode each document into a word-level TF-IDF representation; [^3] Data tokenization is independent of the downstream model. Here, we use the sklearn text-vectorizer tokenizer. we then perform balanced $k$-means clustering on these representations to obtain approximately balanced subsets of the data on which to train each x-elm. Further details on the balanced $k$-means clustering method can be found in Gururangan et al. (2023). This allocation method uses no language information outside of what is inherent in the text (e.g., script, vocabulary).
@@
We also consider segmenting the corpus by language identity. [^4] This requires knowledge of each document’s language. We use the language tags provided with mC4. Rather than balancing the amount of data allocated to each cluster in this setting, we keep the number of languages per cluster fixed. Specifically, we learn a balanced hierarchical clustering of the languages (Figure 2). We build this hierarchy using the language similarity metrics in lang2vec Littell et al. (2017), which represents languages based on linguistic features in resources such as WALS [^5] World Atlas of Language Structures, https://wals.info/ and estimates language similarity with distance in this feature space. We initialize each cluster with a single language; at each step, we merge each cluster with exactly one other based on the minimum distances between cluster centroids. We then group languages by the resulting hierarchy and desired number of experts. When the number of languages equals the number of experts, typological clustering results in monolingual training, as every language is assigned a separate expert.
Figure 2: Hierarchical clustering of languages used to train our x-elm ensembles.
@@ 3.3 Inference with x-elms
We evaluate multiple methods for performing inference with x-elms:
@@
This method performs inference with a single expert chosen prior to evaluation, which incurs the same inference cost as the dense baselines. When evaluating Typology experts on a particular language $\ell$, we choose the expert that included $\ell$ in the set of languages on which they continued pretraining. Similarly, when evaluating TF-IDF, we choose the x-elm trained on the highest percentage of $\ell$’s data.
@@
We also consider ensembling TF-IDF experts by adapting the c-BTM routing method Gururangan et al. (2023). Here, we calculate ensembling $\alpha$, or weights, over the experts at each inference step based on the prior context’s TF-IDF distance from the experts’ $k$-means centroids. These weights are then used to ensemble the output probabilities from each expert.
More specifically, given a probability from each expert $p_{e}(x_{t}\mid x_{<t})$ and the corresponding ensemble weight $\alpha_{e}=p(e\mid x_{<t})\propto\textrm{exp}(-\textrm{dist}(x_{<t},c_{e})^{2}/T)$, the probability of the ensemble is
$p_{E}(x_{t}\mid x_{<t})=\sum_{e\in E}\alpha_{e}\cdot p_{e}(x_{t}\mid x_{<t})$
Here, $\textrm{dist}(x_{<t},c_{e})$ is obtained by embedding $x_{<t}$ with the learned TF-IDF vectorizer and calculating the Euclidean distance from $c_{e}$ (the centroid over the data representations allocated to expert $e$), and $T$ is a temperature parameter over the ensemble weight distribution. Further details and motivation are provided in Gururangan et al. (2023).
Ensembling x-elm outputs increases the cost of inference relative to dense models or top-1 inference. However, it can potentially better fit different subsets of data in a diverse evaluation set. We also do not assume we know the language of each example when ensembling, which makes this approach more flexible than the top-1 setting. In most cases, we ensemble all $k$ experts; however, we can also reduce computational costs by sparsifying the ensemble and only activating the $m$ ($<k$) experts that most contribute to an example: $p_{E}(x_{t}\mid x_{<t})=\sum_{e\in E}\alpha_{e}\cdot p_{e}(x_{t}\mid x_{<t}):\alpha_{e}\in\textrm{top-m}(\bf{\alpha}_{E})$. Appendix Table 8 presents the performance tradeoff with sparser TF-IDF ensembles.
Figure 3: Average and language-specific (EN and SW) perplexities across different expert counts $k$ (Num. Experts). In each evaluation setting, we compare clustering training data for experts with the TF-IDFtop1 (square) and Linguistic Typology (triangle) methods (§3.2). The best choice of $k$ for each setting is marked with a star.
@@ 4 Hierarchical Multi-Round Training
We previously described a single round of training for x-elm3.1). However, BTM can also be used iteratively to train new experts seeded with those learned in a prior round. The multilingual setting provides a natural extension of multi-round training that leverages typological structure.
@@
HMR pretraining gives multiple benefits over single-round BTM. In particular, HMR training saves compute and more easily adapts our x-elms to new settings. A specific application of this is adding new languages to the model: while updating dense multilingual LMs with new languages is difficult and can lead to catastrophic forgetting of existing languages Winata et al. (2023), hierarchically training an expert on a new language adds it to the x-elm set without altering the existing information in other experts. We evaluate HMR training with this use case in §6.3.
@@ 5.1 Pretraining Data and Languages
We train our x-elms on mC4, an open-source, multilingual pretraining corpus derived from CommonCrawl Xue et al. (2021). [^6] While one could also continue pretraining with the same corpus that the seed LM was trained on, the pretraining data for XGLM is not publicly available. mC4 provides language tags for each document in the corpus, which were automatically assigned with cld3777https://github.com/google/cld3 when the dataset was constructed; we use these language tags during typological clustering (§3.2). We focus our experiments on the 16 highest-resourced languages out of the 30 languages on which the seed LM, XGLM-1.7B, was trained. For languages with significantly more data than the others (e.g., English), we subsample their data to the first 1,024 shards. Appendix Table 5 gives the languages and data quantities in our pretraining corpus.
@@ 5.2 Pretraining Settings
5.3 Baselines
We compare the performance of our x-elm experiments to two dense baselines: XGLM, which is the 1.7B parameter seed model used to initialize each expert prior to x-BTM training, and a single dense model, which is also initialized with XGLM-1.7B’s weights and then trained with the full data and compute budget split across experts in other x-elm settings. This is equivalent to $k$=1 experts in cases where we vary the number of experts.
Since the curse of multilinguality is often evaluated in comparison to monolingual modeling, We also consider the setting where we train monolingual expert models on each target language (§6.1). Given that we consider sixteen languages in our x-elm experiments, this corresponds to the $k=16$ typological clustering setting.
@@ 5.4 Perplexity Evaluation
6 Language Modeling Experiments
We now test the effectiveness of sparse language modeling in the multilingual setting. First, we determine the optimal number of clusters for our given compute budget and dataset (§6.1). We then demonstrate that x-elms outperform comparable dense models on seen languages (§6.2) and more effectively adapt to new, unseen languages (§6.3).
6.1 Choosing the Number of x-elms
We first consider which number of experts gives the best multilingual language modeling performance. Figure 3 compares the choice of $k\in\{1,4,8,16\}$ x-elms when trained on 10.5B tokens. [^8] The $k=16$ setting is equivalent to training monolingual experts for every language. Full results are in Table 1 for $k=8$ and Appendix Table 7 for $k=4$ and $k=16$. $k=8$ is the best-performing setting on 75% of languages when clustering with TF-IDF and for 15 of the 16 pretraining languages when clustering by language similarity. Furthermore, typological clustering consistently outperforms TF-IDF.
These experiments indicate that, for the budget we evaluate, the best overall x-elm setting is bilingual models (k=8) clustered by language similarity. This result is surprising, as it is intuitive to assume that simply continuing to pretrain each expert on a single language (i.e., the $k=16$ setting) would lead to better perplexity. We find that one language, Swahili, does benefit from the monolingual $k=16$ setting—possibly because Swahili is paired with a distant language (Vietnamese) by the typological clustering process. However, perplexity is higher in the $k=16$ setting for all other languages, and in some cases, even underperforms the dense ($k=1$) model.